OpenStudy (anonymous):

What is the solution of the equation? 4.7x + 3.8 = 13.2 A. 3.6 B. 3.7 C. 2 D. 3

OpenStudy (anonymous):

In order to solve this problem first start by isolating the x value. So to do so you must first subtract 3.8 from both sides

OpenStudy (anonymous):

4.7x+3.8=13.2 -3.8 -3.8 4.7x=9.4 then divide each side by 4.7 therefor x will =2